# Random pairing MLE for estimation of item parameters in Rasch model

### v1

@@ Appendix C Proof of Lemma 4
    There exist some $\alpha_{1},\alpha_{2}$ obeying $0<\alpha_{1}\leq\alpha_{2}$ such that any $t\in\mathbb{N}$,
    $\|\bm{\theta}^{t}-\widehat{\bm{\theta}}\|_{\bm{L}_{Lz}}\leq(1-\eta\alpha_{1})^{t}\|\bm{\theta}^{0}-\widehat{\bm{\theta}}\|_{\bm{L}_{Lz}},$
@@
    For any $k,l$ and iteration $t\geq 0$,
    $\left|\left(\theta_{k}^{t}-\theta_{l}^{t}\right)-\left(\theta_{k}^{\star}-\theta_{l}^{\star}\right)\right|\leq C_{2}\kappa_{1}\sqrt{\frac{\log(n)}{\lambda_{m-1}(\bm{L}_{L})}}$ (27)
@@
    For any $k,l$ and iteration $t\geq 0$,
    $\left|(\bm{e}_{k}-\bm{e}_{l})^{\top}\bm{L}_{Lz}^{\dagger}\bm{r}^{t}\right|\leq C_{3}\kappa_{1}^{3}\frac{(d_{\max})^{2}\log^{2}(n)}{(\lambda_{m-1}(\bm{L}_{L}))^{3}}$
@@
Then the condition (26) holds as long as $np\geq C_{4}\kappa_{1}^{2}\kappa_{2}\log^{3}(n)$ for some large enough constant $C_{4}$. Invoke Lemma 6 to see that for any $(k,l,t)$,
$\left|(\bm{e}_{k}-\bm{e}_{l})^{\top}\bm{L}_{Lz}^{\dagger}\bm{r}^{t}\right|\leq C_{5}\kappa_{1}^{6}\frac{\log^{2}(n)}{np},$ (28)
@@
$\widehat{\bm{\delta}}$ $=\lim_{t\rightarrow\infty}\bm{\delta}^{t}=-\bm{L}_{Lz}^{\dagger}\bm{B}\widehat{\bm{\epsilon}}+\underbrace{\eta\lim_{t\rightarrow\infty}\sum_{i=0}^{t-1}(1-\eta)^{t-1-i}\bm{L}_{Lz}^{\dagger}\bm{r}^{i}}_{\eqqcolon\bm{r}}.$ (29)
[…]
$|r_{k}-r_{j}|\leq\eta\lim_{t\rightarrow 0}\sum_{i=0}^{t-1}(1-\eta)^{t-1-i}C_{5}\kappa_{1}^{6}\frac{\log^{2}(n)}{np}=C_{5}\kappa_{1}^{6}\frac{\log^{2}(n)}{np}.$
@@
Appendix D Proof of Theorem 2
We start with the proof of (5). By Lemma 4 we can express the MLE estimation error $\widehat{\bm{\theta}}-\bm{\theta}^{\star}$ as
@@ Appendix D Proof of Theorem 2
$\bm{L}_{Lz}^{\dagger}\bm{B}\widehat{\bm{\epsilon}}=\sum_{i>j:(i,j)\in\mathcal{E}_{Y}}\sum_{l=1}^{L_{ij}}\underbrace{\epsilon_{ij}^{(l)}\bm{L}_{Lz}^{\dagger}(\bm{e}_{i}-\bm{e}_{j})}_{\eqqcolon\bm{u}_{ij}^{(l)}}.$
It is easy to see that conditional on $\mathcal{G}_{Y}$, $\{\bm{u}_{ij}^{(l)}\}_{i,j,l}$ is a set of independent zero-mean random variables. Thus we can expand $\mathbb{E}[\|\bm{L}_{Lz}^{\dagger}\bm{B}\widehat{\bm{\epsilon}}\|^{2}]$ to be
$\mathbb{E}\left[\|\bm{L}_{Lz}^{\dagger}\bm{B}\widehat{\bm{\epsilon}}\|^{2}\right]$ $=\mathbb{E}\sum_{i>j:(i,j)\in\mathcal{E}_{Y}}\sum_{l=1}^{L_{ij}}\bm{u}_{ij}^{(l)\top}\bm{u}_{ij}^{(l)}$
$=\mathbb{E}\sum_{i>j:(i,j)\in\mathcal{E}_{Y}}\sum_{l=1}^{L_{ij}}\mathrm{Trace}\left(\bm{u}_{ij}^{(l)}\bm{u}_{ij}^{(l)\top}\right)$
[…]
@@
$z_{ij}=\mathrm{Var}(\epsilon_{ji}^{(l)})=\sigma^{\prime}(\theta_{i}^{\star}-\theta_{j}^{\star})=\frac{e^{\theta_{i}^{\star}}e^{\theta_{j}^{\star}}}{(e^{\theta_{i}^{\star}}+e^{\theta_{j}^{\star}})^{2}},$
@@
Moreover, $\{\epsilon_{ij}^{(l)}\}_{i,j,l}$ is a set of sub-Gaussian random variable with variance proxy $1/z_{ij}$ (see, e.g., Section 2.5 in [Ver18]) and $1/z_{ij}\leq 4\kappa_{1}$ by Lemma 9. Applying Hanson-Wright inequality (see [RV13]), for any scalar $a>0$ we have
@@
[…]
@@
This implies that $\|\bm{L}_{Lz}^{\dagger}\|\leq 16\kappa_{1}\kappa_{2}/(np)$ and $\|\bm{L}_{L\hat{z}}^{\dagger}\|\leq 32\kappa_{1}\kappa_{2}/(np)$. Using the perturbation bound of pseudo-inverse (see Theorem 4.1 in [Wed73]), we have
@@
Appendix E Proofs for inference
In this section, we present the analysis for the inference results we presented in Section 3.3.
E.1 Proof of Theorem 3
We condition the whole analysis on the high probability event when Lemmas 4 and 3 holds. Let $\widehat{\bm{\delta}}\coloneqq\widehat{\bm{\theta}}-\bm{\theta}^{\star}$. Expanding (15), we have
$\widehat{\bm{\delta}}$ $=-\bm{L}_{Lz}^{\dagger}\bm{B}\widehat{\bm{\epsilon}}+\bm{r}$
$=-\bm{L}_{Lz}^{\dagger}\sum_{i>j:(i,j)\in\mathcal{E}_{Y}}\sum_{l=1}^{L_{ij}}\left[Y_{ji}^{(l)}-\sigma(\theta_{i}^{\star}-\theta_{j}^{\star})\right](\bm{e}_{i}-\bm{e}_{j})+\bm{r},$
We start with the case with $\widetilde{\bm{z}}=\bm{z}$. Consider $\bm{L}_{Lz}\widehat{\bm{\delta}}$. As $\bm{L}_{Lz}\bm{1}_{m}=\bm{0}$, $\lambda_{m-1}(\bm{L}_{Lz})>0$, and $\widehat{\bm{\delta}}^{\top}\bm{1}_{m}=0$,
$\bm{L}_{Lz}^{1/2}\widehat{\bm{\delta}}$ $=-(\bm{L}_{Lz}^{\dagger})^{1/2}\bm{B}\widehat{\bm{\epsilon}}+\bm{L}_{Lz}^{1/2}\bm{r}$
$=-\sum_{i>j:(i,j)\in\mathcal{E}_{Y}}\sum_{l=1}^{L_{ij}}\underbrace{\left[Y_{ji}^{(l)}-\sigma(\theta_{i}^{\star}-\theta_{j}^{\star})\right](\bm{L}_{Lz}^{\dagger})^{1/2}(\bm{e}_{i}-\bm{e}_{j})}_{\eqqcolon\bm{u}_{ij}^{(l)}}+\bm{L}_{Lz}^{1/2}\bm{r}.$
Conditional on $(\mathcal{E}_{Y},\{L_{ij}\})$, $\bm{u}_{ij}^{(l)}$ are independent random variables. It is easy to see that $\bm{u}_{ij}^{(l)}$ is zero-mean and has covariance
$\mathbb{E}\left[\bm{u}_{ij}^{(l)}\bm{u}_{ij}^{(l)\top}\right]$ $=z_{ij}(\bm{L}_{Lz}^{\dagger})^{1/2}(\bm{e}_{i}-\bm{e}_{j})(\bm{e}_{i}-\bm{e}_{j})^{\top}(\bm{L}_{Lz}^{\dagger})^{1/2}.$
It is also bounded in third moment of its spectral norm by
$\mathbb{E}\left[\|\bm{u}_{ij}^{(l)}\|^{3}\right]\leq 2^{3/2}\|\bm{L}_{Lz}^{\dagger}\|^{3/2}\leq\frac{2^{15/2}\kappa_{1}^{3/2}\kappa_{2}^{3/2}}{(np)^{3/2}},$
$\sum_{i>j:(i,j)\in\mathcal{E}_{Y}}\sum_{l=1}^{L_{ij}}\mathbb{E}\left[\bm{u}_{ij}^{(l)}\bm{u}_{ij}^{(l)\top}\right]$ $=(\bm{L}_{Lz}^{\dagger})^{1/2}\sum_{i>j:(i,j)\in\mathcal{E}_{Y}}L_{ij}z_{ij}(\bm{e}_{i}-\bm{e}_{j})(\bm{e}_{i}-\bm{e}_{j})^{\top}(\bm{L}_{Lz}^{\dagger})^{1/2}$
$=(\bm{L}_{Lz}^{\dagger})^{1/2}\bm{L}_{Lz}(\bm{L}_{Lz}^{\dagger})^{1/2}$
$=\bm{I}_{m}-\frac{1}{m}\bm{1}_{m}\bm{1}_{m}^{\top}.$
The last line holds since $\bm{L}_{Lz}\bm{1}_{m}=\bm{0}$ and $\lambda_{m-1}(\bm{L}_{Lz})>0$. Now using multivariate CLT (see, e.g., [Rai19]), we have that as $\sum_{i>j:(i,j)\in\mathcal{E}_{Y}}L_{ij}\rightarrow\infty$,

### v2

@@ C.1 Proof of Theorem 2
    There exist some $\alpha_{1},\alpha_{2}$ obeying $0<\alpha_{1}\leq\alpha_{2}$ such that any $\tau\in\mathbb{N}$,
    $\|\bm{\theta}^{\tau}-\widehat{\bm{\theta}}\|_{\bm{L}_{Lz}}\leq(1-\eta\alpha_{1})^{\tau}\|\bm{\theta}^{0}-\widehat{\bm{\theta}}\|_{\bm{L}_{Lz}},$
@@
    For any $k,l$ and iteration $\tau\geq 0$,
    $\left|\left(\theta_{k}^{\tau}-\theta_{l}^{\tau}\right)-\left(\theta_{k}^{\star}-\theta_{l}^{\star}\right)\right|\leq C_{2}\kappa_{1}\sqrt{\frac{\log(n)}{\lambda_{m-1}(\bm{L}_{L})}}$ (31)
@@
    For any $k,l$ and iteration $\tau\geq 0$,
    $\left|(\bm{e}_{k}-\bm{e}_{l})^{\top}\bm{L}_{Lz}^{\dagger}\bm{r}^{\tau}\right|\leq C_{3}\kappa_{1}^{3}\frac{(d_{\max})^{2}\log^{2}(n)}{(\lambda_{m-1}(\bm{L}_{L}))^{3}}$
@@
Then the condition (30) holds as long as $np\geq C_{4}\kappa_{1}^{2}\kappa_{2}\log^{3}(n)$ for some large enough constant $C_{4}$. Invoke Lemma 5 to see that for any $(k,l,\tau)$,
$\left|(\bm{e}_{k}-\bm{e}_{l})^{\top}\bm{L}_{Lz}^{\dagger}\bm{r}^{\tau}\right|\leq C_{5}\kappa_{1}^{6}\frac{\log^{2}(n)}{np},$ (32)
@@
$\widehat{\bm{\delta}}$ $=\lim_{\tau\rightarrow\infty}\bm{\delta}^{\tau}=-\bm{L}_{Lz}^{\dagger}\bm{B}\widehat{\bm{\epsilon}}+\underbrace{\eta\lim_{\tau\rightarrow\infty}\sum_{i=0}^{\tau-1}(1-\eta)^{\tau-1-i}\bm{L}_{Lz}^{\dagger}\bm{r}^{i}}_{\eqqcolon\bm{r}}.$ (33)
It remains to control the $\ell_{\infty}$ norm of $\bm{r}$. For any $(k,l$), (32) shows that
$|r_{k}-r_{j}|\leq\eta\lim_{\tau\rightarrow 0}\sum_{i=0}^{\tau-1}(1-\eta)^{\tau-1-i}C_{5}\kappa_{1}^{6}\frac{\log^{2}(n)}{np}=C_{5}\kappa_{1}^{6}\frac{\log^{2}(n)}{np}.$
@@
C.1.1 Proof of Lemma 5
Lemma 5 is a direct combination of Lemmas 7, Lemma 8, and the proof of Lemma 6 in [47]. The results therein describe controls the error for the same dynamic with a more general setting. Thus it is directly applicable to our case. For clarity, in this section we explain the connection and relate our notations with the ones used in [47].
For any $(k,l)\in[m]^{2}$, let
$Q_{kl}\coloneqq C_{1}\kappa_{1}^{3}\frac{(d_{\max})^{2}\log^{2}(n)}{(\lambda_{m-1}(\bm{L}_{Lz}))^{3}}\qquad\text{and}\qquad B_{kl}\coloneqq C_{2}\kappa_{1}\sqrt{\frac{\log(n)}{\lambda_{m-1}(\bm{L}_{Lz})}},$
where $C_{1},C_{2}$ are some constants. Note that our setting in this paper is unweighted and has different number of observation for each edge in $\mathcal{G}_{Y}$. Moreover $\bm{L}_{wz}$ in [47] correspond to $\bm{L}_{Lz}$ in this paper. Then Lemma 3 and Lemma 4 in [47] imply that $Q_{kl}$ and $B_{kl}$ satisfies equation (9) therein. Furthermore, as we have assumed in(30), as long as $C_{2}$ is large enough, $Q_{kl}\leq 4B_{kl}$ for any $(k,l)$. Thus we can invoke Lemma 2 as well as all its proof. Lemma 20 in [47] implies the first two items in Lemma 5 herein. Lemma 19 in [47] implies the item 3. Finally, item 4 appears in the second-to-last equation block in the proof of Lemma 19 in [47].
C.2 Proof of Proposition 3
We condition the whole analysis on the high probability event when Lemmas 2 and 3 hold, which happens with probability at least $1-O(n^{-10})$.
We can expand $\bm{L}_{Lz}^{\dagger}\bm{B}\widehat{\bm{\epsilon}}$ as
$\bm{L}_{Lz}^{\dagger}\bm{B}\widehat{\bm{\epsilon}}=-\bm{L}_{Lz}^{\dagger}\sum_{i>j:(i,j)\in\mathcal{E}_{Y}}\sum_{t:L_{ij}^{t}=1}\left[Y_{ji}^{t}-\sigma(\theta_{i}^{\star}-\theta_{j}^{\star})\right](\bm{e}_{i}-\bm{e}_{j})+\bm{r}$
Consider $\bm{L}_{Lz}(\widehat{\bm{\theta}}-\bm{\theta}^{\star})$. As $\bm{L}_{Lz}\bm{1}_{m}=\bm{0}$, $\lambda_{m-1}(\bm{L}_{Lz})>0$, and $\widehat{\bm{\delta}}^{\top}\bm{1}_{m}=0$,
$(\widehat{\bm{\theta}}-\bm{\theta}^{\star})$ $=-\bm{L}_{Lz}^{\dagger}\bm{B}\widehat{\bm{\epsilon}}+\bm{r}$
$=-\bm{L}_{Lz}^{\dagger}\sum_{i>j:(i,j)\in\mathcal{E}_{Y}}\sum_{t:L_{ij}^{t}=1}\underbrace{\left[Y_{ji}^{t}-\sigma(\theta_{i}^{\star}-\theta_{j}^{\star})\right](\bm{e}_{i}-\bm{e}_{j})}_{\eqqcolon\bm{u}_{ij}^{t}}+\bm{r}.$
Conditional on $(\mathcal{E}_{Y},\{t:L_{ij}^{t}=1\})$, $\bm{u}_{ij}^{t}$ are independent random variables. It is also easy to see that $\bm{u}_{ij}^{t}$ is zero-mean and has covariance
$\mathbb{E}\left[\bm{u}_{ij}^{t}\bm{u}_{ij}^{t\top}\right]$ $=z_{ij}(\bm{e}_{i}-\bm{e}_{j})(\bm{e}_{i}-\bm{e}_{j})^{\top}.$
When rescaled by $(\bm{L}_{Lz}^{\dagger})^{1/2}$, it is also bounded in the third moment of spectral norm by
$\mathbb{E}\left[\|(\bm{L}_{Lz}^{\dagger})^{1/2}\bm{u}_{ij}^{t}\|^{3}\right]\leq 2^{3/2}\|\bm{L}_{Lz}^{\dagger}\|^{3/2}\leq\frac{2^{15/2}\kappa_{1}^{3/2}\kappa_{2}^{3/2}}{(np)^{3/2}},$ (34)
$\sum_{i>j:(i,j)\in\mathcal{E}_{Y}}\sum_{t:L_{ij}^{t}=1}\mathbb{E}\left[\bm{u}_{ij}^{t}\bm{u}_{ij}^{t\top}\right]$ $=\sum_{i>j:(i,j)\in\mathcal{E}_{Y}}L_{ij}z_{ij}(\bm{e}_{i}-\bm{e}_{j})(\bm{e}_{i}-\bm{e}_{j})^{\top}$
$=\bm{L}_{Lz}$
The last line holds since $\bm{L}_{Lz}\bm{1}_{m}=\bm{0}$ and $\lambda_{m-1}(\bm{L}_{Lz})>0$. Now using multivariate Berry–Esseen theorem (see, e.g., [36]), let $\widetilde{\bm{x}}$ be a random variable such that
$\bm{x}\sim\mathcal{N}\left(\bm{0}_{m},\bm{L}_{Lz}^{\dagger}\right),$
then
$\sup_{A\in\mathcal{C}_{m}}\left|\mathbb{P}\left[-\bm{L}_{Lz}^{\dagger}\bm{B}\widehat{\bm{\epsilon}}\in A\mid\mathcal{G}_{Y}\right]-\mathbb{P}(\bm{x}\in A)\right|$ $\leq C_{2}m^{1/4}\sum_{i>j:(i,j)\in\mathcal{E}_{Y}}\sum_{t:L_{ij}^{t}=1}\mathbb{E}\left[\|(\bm{L}_{Lz}^{\dagger})^{1/2}\bm{u}_{ij}^{t}\|^{3}\right]$
$\overset{\text{(i)}}{\leq}C_{3}\frac{m^{5/4}\kappa_{1}^{3/2}\kappa_{2}^{3/2}}{(np)^{1/2}}.$
Here (i) uses 2 and (34), and $C_{2},C_{3}$ are some absolute constants. We then reach the desired conclusion by adding the probability upper bound that Lemmas 2 and 3 fail.
C.3 Proof of Proposition 4
We start with the proof of (6). By Theorem 2 we can express the MLE estimation error $\widehat{\bm{\theta}}-\bm{\theta}^{\star}$ as
@@ C.3 Proof of Proposition 4
$\bm{L}_{Lz}^{\dagger}\bm{B}\widehat{\bm{\epsilon}}=\sum_{i>j:(i,j)\in\mathcal{E}_{Y}}\sum_{t:L_{ij}^{t}=1}\underbrace{\epsilon_{ij}^{t}\bm{L}_{Lz}^{\dagger}(\bm{e}_{i}-\bm{e}_{j})}_{\eqqcolon\bm{u}_{ij}^{t}}.$
It is easy to see that conditional on $(\mathcal{E}_{Y},\{t:L_{ij}^{t}=1\})$, $\{\bm{u}_{ij}^{t}\}_{i,j,t}$ is a set of independent zero-mean random variables. Thus we can expand $\mathbb{E}[\|\bm{L}_{Lz}^{\dagger}\bm{B}\widehat{\bm{\epsilon}}\|^{2}]$ to be
$\mathbb{E}\left[\|\bm{L}_{Lz}^{\dagger}\bm{B}\widehat{\bm{\epsilon}}\|^{2}\right]$ $=\mathbb{E}\sum_{i>j:(i,j)\in\mathcal{E}_{Y}}\sum_{t:L_{ij}^{t}=1}\bm{u}_{ij}^{t\top}\bm{u}_{ij}^{t}$
$=\mathbb{E}\sum_{i>j:(i,j)\in\mathcal{E}_{Y}}\sum_{t:L_{ij}^{t}=1}\mathrm{Trace}\left(\bm{u}_{ij}^{t}\bm{u}_{ij}^{t\top}\right)$
$\overset{\text{(i)}}{=}\mathrm{Trace}\left(\sum_{i>j:(i,j)\in\mathcal{E}_{Y}}\sum_{t:L_{ij}^{t}=1}\bm{L}_{Lz}^{\dagger}z_{ij}(\bm{e}_{i}-\bm{e}_{j})(\bm{e}_{i}-\bm{e}_{j})^{\top}\bm{L}_{Lz}^{\dagger}\right)$
$\overset{\text{(ii)}}{=}\mathrm{Trace}(\bm{L}_{Lz}^{\dagger}).$ (36)
@@
$z_{ij}=\mathrm{Var}(\epsilon_{ji}^{t})=\sigma^{\prime}(\theta_{i}^{\star}-\theta_{j}^{\star})=\frac{e^{\theta_{i}^{\star}}e^{\theta_{j}^{\star}}}{(e^{\theta_{i}^{\star}}+e^{\theta_{j}^{\star}})^{2}},$
@@
Moreover, $\{\epsilon_{ij}^{t}\}_{(i,j,t):i>j,(i,j)\in\mathcal{E}_{Y},L_{ij}^{t}=1}$ is a set of sub-Gaussian random variable with variance proxy $1/z_{ij}$ (see, e.g., Section 2.5 in [44] for the definition of sub-Gaussian random variable) and $1/z_{ij}\leq 4\kappa_{1}$ by Lemma 11. Applying Hanson-Wright inequality (see [39]), for any scalar $a>0$ we have
@@
$\overset{\text{(i)}}{=}\sqrt{\mathrm{Trace}\left(\bm{L}_{Lz}^{\dagger}\bm{L}_{Lz}^{\dagger}\right)}$
$\overset{\text{(ii)}}{\leq}\frac{\sqrt{m}}{\lambda_{m-1}(\bm{L}_{Lz})}\overset{\text{(iii)}}{\leq}\frac{16\kappa_{1}\kappa_{2}\sqrt{m}}{np}.$
@@
This implies that $\|\bm{L}_{Lz}^{\dagger}\|\leq 16\kappa_{1}\kappa_{2}/(np)$ and $\|\bm{L}_{L\hat{z}}^{\dagger}\|\leq 32\kappa_{1}\kappa_{2}/(np)$. Using the perturbation bound of pseudo-inverse (see Theorem 4.1 in [46]), we have
@@
Appendix D Proofs for Section 3.3
In this section we prove the results for Section 3.3, including Theorem 3, Theorem 4, Proposition 5, and (9). To facilitate our analysis, we decompose the loss function by each user and random split. For $\mathrm{WP\text{-}MLE}$, recall that we let the loss function associated with user $t$ to be
$\mathcal{L}_{\mathrm{WP}}^{(t)}(\bm{\theta})\coloneqq-\sum_{\begin{subarray}{c}(i,j):i>j\\ (t,i),(t,j)\in\mathcal{G}_{X}\end{subarray}}\frac{\widetilde{m}_{t}}{m_{t}(m_{t}-1)}\left[\log\left(\frac{e^{\theta_{i}}}{e^{\theta_{i}}+e^{\theta_{j}}}\right)\mathds{1}\{X_{ti}>X_{tj}\}\right.\\ \left.\hskip 80.00012pt{}+\log\left(\frac{e^{\theta_{j}}}{e^{\theta_{i}}+e^{\theta_{j}}}\right)\mathds{1}\{X_{ti}<X_{tj}\}\right].$
We have also defined the loss function associated with random split $k$ and user $t$ for $\mathrm{MRP\text{-}MLE}$ as
$\mathcal{L}_{k}^{(t)}(\bm{\theta})=-\sum_{\begin{subarray}{c}(i,j):i>j,\\ (i,j,t)\in\Omega_{k}\end{subarray}}\left[\log\left(\frac{e^{\theta_{i}}}{e^{\theta_{i}}+e^{\theta_{j}}}\right)\mathds{1}\{X_{ti}>X_{tj}\}+\log\left(\frac{e^{\theta_{j}}}{e^{\theta_{i}}+e^{\theta_{j}}}\right)\mathds{1}\{X_{ti}<X_{tj}\}\right],$
where $\Omega_{k}$ denotes the set of paired item-item-user tuples for the $k$-th split.
D.1 Proof of Theorem 3
We first fix a random split $k$. By mean value theorem